You are a Grade 10 representative on the student council and you have been assigned the task of finding a location to House of dance for students of CISE you have narrowed it down to three options which are outlined below.


ancient abyss is a location that charges a flat fee of $200 and $4 dollars per person, beautiful babylon is a location that charges a $99 flat fee and $6 per person, crazy Cleoparta is a location that charges $280 flat fee and $3 per person you will analyze the three options and put together a proposal to submit to the student council outlining your recommendations


1
Expert's answer
2021-03-02T01:05:09-0500

Let "n=" the number of students.

Ancient abyss: "C_A(n)=200+4n"

Beautiful babylon: "C_B(n)=99+6n"  

Crazy Cleoparta: "C_C(n)=280+3n"  


"200+4n=99+6n=>n=50.5"

For "n\le50"


"C_B<C_A<C_C"

"280+3n=99+6n=>n=60\dfrac{1}{3}"

For "51\leq n\le60"


"C_A<C_B<C_C"


"200+4n=280+3n=>n=80"

For "51\leq n\leq79"


"C_A<C_C<C_B"

For "n=80"


"C_A=C_C<C_B"

For "n>80"


"C_C<C_A<C_B"